OpenStudy (anonymous):

Oil spilled from a ruptured tanker is spreading in a circular pattern. If the radius of the circle is increasing at the constant rate of 1.5 feet per second, how fast is the area of the spill increasing when the radius of the spill is 42 feet?

OpenStudy (goformit100):

use speed time formula...

OpenStudy (anonymous):

I've never heard of that.

OpenStudy (anonymous):

related rates right?

OpenStudy (anonymous):

Yeah, related rates.

OpenStudy (anonymous):

\(A=\pi r^2\) and so \(A'=2\pi rr'\)

OpenStudy (anonymous):

you are told \(r'=1.5\) and \(r=42\) plug in the numbers and get \(A'\) in one step

OpenStudy (anonymous):

wow, that was easy. I'm terrible at these.

OpenStudy (goformit100):

@bloodrain

OpenStudy (anonymous):

yeah, i have to admit that one was real easy they get harder

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

well only the formulas get harder to figure out, i am willing to bet that you knew the formula for the area of a circle

OpenStudy (anonymous):

yeah, I was just thinking like A=pi(r^t/1.5)^2 then I was super confused

OpenStudy (anonymous):

That made absolutely no sense to me.

OpenStudy (anonymous):

The minute hand on a watch is 8mm long and the hour hand is 4 mm long. How fast is the distance between the tips of the hands changing at one o'clock? THat's the hardest one I have done myself.